Given the functions: ATC= x-6+(30/x) and AR =10+(15/x), where x represents output, find TC, TR, and the output level when net revenue is zero.

1
Expert's answer
2021-02-18T15:36:25-0500

"\bold {Answers}"

When net revenue is zero, output = 1 unit or 15 units.

When output = 1 unit, TR = TC = $25, and,

When output = 15 units, TR = TC = $165

"\bold {Solutions}"

We are given:


"ATC = x - 6 + \dfrac {30}{x}" and "AR = 10 + \dfrac {15}{x}"


Now,

"TC = x \u00d7 ATC"

"= x (x - 6 + \dfrac {30}{x})"

"= x^2 - 6x + 30"


"TR = x \u00d7 AR"

"= x (10 + \dfrac {15}{x})"

"= 10x + 15"


When net revenue is zero:

"TR - TC = 0"

"=> x^2 - 6x + 30 -( 10x + 15 )= 0"

"=> x^2 -16x + 15 = 0"

"=> x^2 - x - 15x + 15 = 0"

"=> x(x-1)-15(x-1) = 0"

"=> (x-1)(x-15) = 0"

"\therefore \space x = 1 \space or \space x = 15"


Thus, when net revenue is zero, output is either 1 unit or 15 units.


When "x = 1"

"TR = TC = (1)^2 - 6(1) + 30"

"= 1 - 6 + 30"

"\bold {= \$25}"


When "x = 15"

"TC = TR = 10(15) + 15"

"= 150 + 15"

"\bold {= \$165}"